# Webs Generated by Products of convex and homogeneous Foliations on $\mathbb{P}^{2}$

### v1

@@ 2. Convex Reduced Foliations
Proof.:  If $\operatorname{Tang}(\mathcal{F}_{l},\mathcal{F}_{d})=l_{1}\cup\cdots\cup l_{s}$ where $l_{j}$ are distinct invariant lines of $\mathcal{F}_{l}$ and $\mathcal{F}_{d}$, we know that $\mathcal{F}_{l}$ and $\mathcal{F}_{d}$ have $3l$ and $3d$ invariant lines, respectively, being them
@@
Now we consider product of foliations and one invariant line.
@@
ii. If $s\in(\Sigma_{\mathcal{F}_{1}}^{\text{rad}}\cap\Sigma_{\mathcal{F}_{2}}^{\text{rad}})\setminus l$ with radiality orders $\nu_{1}$ and $\nu_{2}$ respectively. Since $s\not\in l$ we obtain that the direction of $\operatorname{Leg}l$ on the general point of $\check{s}$ does not coincide with any other of $\operatorname{Leg}\mathcal{F}_{1}$ and $\operatorname{Leg}\mathcal{F}_{2}$. By lemma 1.3 we can write
$\operatorname{Leg}\mathcal{W}=(\mathcal{W}_{\nu_{1}+\nu_{2}})\boxtimes(\operatorname{Leg}l\boxtimes\mathcal{W}_{d_{1}+d_{2}-\nu_{1}-\nu_{2}})$
and apply [2, Proposition 2.6] to conclude that $\operatorname{Leg}\mathcal{W}$ is flat.
iii. If $s\in l\setminus\Sigma_{\mathcal{F}_{2}}^{\text{rad}}$ with $\nu_{1}$ being the radiality order of $s$ in $\mathcal{F}_{1}$, then we write $\operatorname{Leg}\mathcal{F}_{1}=\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{d_{1}-\nu_{1}}$. Note that the directions of $\operatorname{Leg}\mathcal{F}_{2}$ and $\mathcal{W}_{d_{1}-\nu_{1}}$ do not coincide at generic $\check{t}\in\check{s}$, and $\operatorname{Tang}(\mathcal{F}_{2},t,s)\leq 1$ (because $s\not\in\Sigma_{\mathcal{F}_{2}}^{\text{rad}}$). Then,
@@
    with $\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-1}$ smooth and transversal to $\check{s}$ and $\mathcal{W}_{1}$ a foliation tangent to $\check{s}$. If $\nu_{1}\geq 2$ the flatness of $\operatorname{Leg}\mathcal{W}$ will follow from [2, Remark 3.10]. On other hand, if $\nu_{1}=1$ we conclude as before from Lemma 1.3 and [2, Proposition 3.9].
iv. If $s\in\Sigma_{\mathcal{F}_{1}}^{\text{rad}}\cap\Sigma_{\mathcal{F}_{2}}^{\text{rad}}\cap l$ with $\nu_{1}$ and $\nu_{2}$ being the orders of $s$, respectively, in $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, then by Lemma 1.3 we can write
$\operatorname{Leg}\mathcal{W}=(\operatorname{Leg}l\boxtimes\mathcal{W}_{\nu_{1}+\nu_{2}})\boxtimes(\mathcal{W}_{d_{1}+d_{2}-\nu_{1}-\nu_{2}})$
with $\mathcal{W}_{d_{1}+d_{2}-\nu_{1}-\nu_{2}}$ smooth and transversal to $\check{s}$. We conclude again by [2, Prop 3.9].
Corollary 2.6.
Let $\mathcal{F}_{1}$, $\mathcal{F}_{2}$ be convex reduced foliations on $\mathbb{P}_{\mathbb{C}}^{2}$ of respective degrees $d_{1}$ and $d_{2}$, such that, $\operatorname{Tang}(\mathcal{F}_{1},\mathcal{F}_{2})$ is a union of invariant lines. For invariant lines $l_{1}$, $l_{2}$ consider the web $\mathcal{W}=l_{1}\boxtimes l_{2}\boxtimes\mathcal{F}_{1}\boxtimes\mathcal{F}_{2}$. Then $\operatorname{Leg}\mathcal{W}=\operatorname{Leg}l_{1}\boxtimes\operatorname{Leg}l_{2}\boxtimes\operatorname{Leg}\mathcal{F}_{1}\boxtimes\operatorname{Leg}\mathcal{F}_{2}$ is flat.
Proof.:  We write in a generic point, $\mathcal{W}_{1}=\operatorname{Leg}\mathcal{F}_{1}=\mathcal{G}_{1}\boxtimes\cdots\boxtimes\mathcal{G}_{n}$ and $\mathcal{W}_{2}=\operatorname{Leg}\mathcal{F}_{2}=\mathcal{H}_{1}\boxtimes\cdots\boxtimes\mathcal{H}_{m}$ where $\mathcal{G}_{i}$ and $\mathcal{H}_{j}$ are foliations for each $i=1,\cdots,n$ and $j=1\cdots,m$. Then,
$K(\operatorname{Leg}\mathcal{W})$ $=$ $\sum_{i=1}^{n}K(\operatorname{Leg}l_{1}\boxtimes\operatorname{Leg}l_{2}\boxtimes\mathcal{G}_{i})+\sum_{j=1}^{m}K(\operatorname{Leg}l_{1}\boxtimes\operatorname{Leg}l_{2}\boxtimes\mathcal{H}_{j})$
$+\sum_{i<j}K(\operatorname{Leg}l_{1}\boxtimes\mathcal{G}_{i}\boxtimes\mathcal{G}_{j})+\sum_{i<j}K(\operatorname{Leg}l_{2}\boxtimes\mathcal{G}_{i}\boxtimes\mathcal{G}_{j})+\sum_{i<j}K(\operatorname{Leg}l_{1}\boxtimes\mathcal{H}_{i}\boxtimes\mathcal{H}_{j})$
$+\sum_{i<j}K(\operatorname{Leg}l_{2}\boxtimes\mathcal{H}_{i}\boxtimes\mathcal{H}_{j})+\sum_{i=1,j=1}^{n,m}K(\operatorname{Leg}l_{1}\boxtimes\mathcal{G}_{i}\boxtimes\mathcal{H}_{j})$
$+\sum_{i=1,j=1}^{n,m}K(\operatorname{Leg}l_{2}\boxtimes\mathcal{G}_{i}\boxtimes\mathcal{H}_{j})+K(\mathcal{W}_{1}\boxtimes\mathcal{W}_{2})$
On the other hand we have
$K(\operatorname{Leg}l\boxtimes\mathcal{W}_{1})=\sum_{i<j}K(\operatorname{Leg}l\boxtimes\mathcal{G}_{i}\boxtimes\mathcal{G}_{j})+K(\mathcal{W}_{1})$
$K(\operatorname{Leg}l_{1}\boxtimes\operatorname{Leg}l_{2}\boxtimes\mathcal{W}_{1})$ $=$ $\sum_{i=1}^{n}K(\operatorname{Leg}l_{1}\boxtimes\operatorname{Leg}l_{2}\boxtimes\mathcal{G}_{i})+\sum_{i<j}K(\operatorname{Leg}l_{1}\boxtimes\mathcal{G}_{i}\boxtimes\mathcal{G}_{j})$
$+\sum_{i<j}K(\operatorname{Leg}l_{2}\boxtimes\mathcal{G}_{i}\boxtimes\mathcal{G}_{j})+K(\mathcal{W}_{1})$
$K(\operatorname{Leg}l\boxtimes\mathcal{W}_{1}\boxtimes\mathcal{W}_{2})$ $=$ $\sum_{i<j}K(\operatorname{Leg}l\boxtimes\mathcal{G}_{i}\boxtimes\mathcal{G}_{j})+\sum_{i<j}K(\operatorname{Leg}l\boxtimes\mathcal{H}_{i}\boxtimes\mathcal{H}_{j})$
$+\sum_{i=1,j=1}^{n,m}K(\operatorname{Leg}l\boxtimes\mathcal{G}_{i}\boxtimes\mathcal{H}_{j})+K(\mathcal{W}_{1}\boxtimes\mathcal{W}_{2})$
Putting all these together with the fact that $K(\mathcal{W}_{i})=K(\mathcal{W}_{1}\boxtimes\mathcal{W}_{2})=K(\operatorname{Leg}(l_{i}\boxtimes\mathcal{F}_{1}\boxtimes\mathcal{F}_{2}))=0$, and also $K(\operatorname{Leg}(l_{1}\boxtimes l_{2}\boxtimes\mathcal{F}_{i}))=0$ by [1, Theorem 1], we obtain $K(\operatorname{Leg}\mathcal{W})=0$.
@@
By [1, Lemma 2.1], we have that
$K(\mathcal{W}^{\prime}\boxtimes\mathcal{W}^{\prime\prime})=K(\mathcal{W}^{\prime})-(n-2)\sum_{i=1}^{n}K(\operatorname{Leg}l_{i}\boxtimes\mathcal{W}^{\prime\prime})+\sum_{1\leq i<j\leq n}K(\operatorname{Leg}l_{i}\boxtimes\operatorname{Leg}l_{j}\boxtimes\mathcal{W}^{\prime\prime})+\left(\begin{matrix}n-1\\ 2\end{matrix}\right)K(\mathcal{W}^{\prime\prime})$
where $\mathcal{W}^{\prime}=\operatorname{Leg}l_{i}\boxtimes\cdots\boxtimes\operatorname{Leg}l_{n}$ and $\mathcal{W}^{\prime\prime}=\operatorname{Leg}\mathcal{F}_{1}\boxtimes\operatorname{Leg}\mathcal{F}_{2}$.
As $K(\mathcal{W}^{\prime})=0$ and by Propositions 2.2, 2.5, and Corollary 2.6, we have that $K(\operatorname{Leg}\mathcal{W})=0$.
Remark 2.7.
Although we can look a statement with $k$ convex reduced foliations, the Proposition 2.4 shows that it is not possible to increase the number of foliations by repeating the argument of Example 2.3 with Fermat Foliations. On the other hand, to the best of our knowledge, starting from degree $8$ there is no other know convex reduced foliations besides Fermat foliation, see for example [6, Problem 9.1].
@@ 3. Homogeneous Foliations
where $d=d_{1}+\cdots+d_{n}$. So we have that $\operatorname{Leg}\mathcal{W}$ is invariant by the dual maps $\check{h}_{\lambda}(a,b)=(\frac{a}{\lambda},\frac{b}{\lambda})$ and so
@@
Let $\mathcal{H}_{1}$, $\mathcal{H}_{2}$ be convex homogeneous foliations on $\mathbb{P}^{2}$ such that, $\operatorname{Tang}(\mathcal{H}_{1},\mathcal{H}_{2})=L_{\infty}\cup l_{1}\cup\cdots\cup l_{k}$ where $l_{j}$ is an invariant line $\forall j=1,\cdots,k$. Consider the web $\mathcal{W}=\mathcal{H}_{1}\boxtimes\mathcal{H}_{2}$. Then $\operatorname{Leg}\mathcal{W}=\operatorname{Leg}\mathcal{H}_{1}\boxtimes\operatorname{Leg}\mathcal{H}_{2}$ is flat.
@@
If $s\in(\Sigma_{\mathcal{H}_{1}}^{rad}\cup\Sigma_{\mathcal{H}_{2}}^{rad})$, we do a similar analysis to that done in the Theorem 2.2 and obtain that $K(\operatorname{Leg}\mathcal{W})$ is holomorphic over $\check{s}$. Thus $K(\operatorname{Leg}\mathcal{W})$ is holomorphic on ˇ$\check{\mathbb{P}}^{2}_{\mathbb{C}}\setminus\check{O}$ and by Lemma 3.1 we have that $\operatorname{Leg}\mathcal{W}$ is flat.

### v2

@@ 2. Convex Reduced Foliations
Proof.:  If $\operatorname{Tang}(\mathcal{F}_{l},\mathcal{F}_{d})=l_{1}\cup\ldots\cup l_{s}$ where $l_{j}$ are distinct invariant lines of $\mathcal{F}_{l}$ and $\mathcal{F}_{d}$, we know that $\mathcal{F}_{l}$ and $\mathcal{F}_{d}$ have $3l$ and $3d$ invariant lines, respectively, being them
@@
In a similar way to Proposition 2.2 we can prove the result with three foliations.
Proposition 2.5.
Let $\mathcal{F}_{1}$, $\mathcal{F}_{2}$, $\mathcal{F}_{3}$ be convex reduced foliations on $\mathbb{P}_{\mathbb{C}}^{2}$ of respective degrees $d_{1}$, $d_{2}$ and $d_{3}$, such that, $\operatorname{Tang}(\mathcal{F}_{i},\mathcal{F}_{j})=C_{ij}^{inv}$ where $\forall i,j=1,2,3$, $C_{ij}^{inv}$ is a product of invariant lines. Consider the web $\mathcal{W}=\mathcal{F}_{1}\boxtimes\mathcal{F}_{2}\boxtimes\mathcal{F}_{3}$. Then $\operatorname{Leg}\mathcal{W}=\operatorname{Leg}\mathcal{F}_{1}\boxtimes\operatorname{Leg}\mathcal{F}_{2}\boxtimes\operatorname{Leg}\mathcal{F}_{3}$ is flat.
$\Delta(\operatorname{Leg}\mathcal{W})=\check{\Sigma}_{\mathcal{F}_{1}}^{rad}\cup\check{\Sigma}_{\mathcal{F}_{2}}^{rad}\cup\check{\Sigma}_{\mathcal{F}_{3}}^{rad}$
Take $s\in\Sigma_{\mathcal{F}_{1}}^{rad}\setminus(\Sigma_{\mathcal{F}_{2}}^{rad}\cup\Sigma_{\mathcal{F}_{3}}^{rad})$, we have that $s$ is a radial singularity of order $\nu_{1}-1$ of $\mathcal{F}_{1}$ and given a generic line $l$ (particularly, not invariant) on $\check{s}$, we can write $\operatorname{Leg}\mathcal{F}_{1}=\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{d_{1}-\nu_{1}}$ with $\mathcal{W}_{\nu_{1}}$ leaving $\check{s}$ invariant and with $\Delta(\mathcal{W}_{\nu_{1}})$ having minimal multiplicity along $\check{s}$ and the $(d_{1}-\nu_{1})$-web $\mathcal{W}_{d_{1}-\nu_{1}}$ is transverse to $\check{s}$ As $s\not\in\Sigma_{\mathcal{F}_{2}}^{rad}\cup\Sigma_{\mathcal{F}_{3}}^{rad}$ we have that $\operatorname{Tang}(\mathcal{F}_{2},l,s)$ and $\operatorname{Tang}(\mathcal{F}_{3},l,s)$ are less or equal than $1$.
  • If $s\not\in(\operatorname{Sing}\mathcal{F}_{2}\cup\operatorname{Sing}\mathcal{F}_{3})$, we can assume $l\cap(\operatorname{Sing}\mathcal{F}_{2}\cup\operatorname{Sing}\mathcal{F}_{3})=\emptyset$, then $\operatorname{Leg}\mathcal{F}_{2}$ and $\operatorname{Leg}\mathcal{F}_{3}$ are completely transversal to $\check{s}$. On the other hand, for example, if one direction of $\mathcal{W}_{d_{1}-\nu_{1}}$ and $\operatorname{Leg}\mathcal{F}_{2}$ coincide, this direction would correspond to a point of $\operatorname{Tang}(\mathcal{F}_{1},\mathcal{F}_{2})=l_{1}\cup\ldots\cup l_{k}$ therefore, we would have $l=l_{j}$ for some $j$, which does not occur. Thus, $\operatorname{Leg}\mathcal{F}_{2}$ and $\operatorname{Leg}\mathcal{F}_{3}$ are transverse to $\mathcal{W}_{d-\nu_{1}}$ and in the same way they are mutually transversal, thus in the neighborhood of $l$ we can write $\operatorname{Leg}\mathcal{W}=\mathcal{W}_{\nu_{1}}\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\operatorname{Leg}\mathcal{F}_{2}\boxtimes\operatorname{Leg}\mathcal{F}_{3})$ with $\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\operatorname{Leg}\mathcal{F}_{2}\boxtimes\operatorname{Leg}\mathcal{F}_{3}$ smooth and transverse to $\check{s}$. Therefore, by [6, Proposition 2.6], the curvature of $\operatorname{Leg}\mathcal{W}$ is holomorphic along $\check{s}$.
  • If $s\in(\operatorname{Sing}\mathcal{F}_{2}\setminus\operatorname{Sing}\mathcal{F}_{3})$ we have $\operatorname{Tang}(\mathcal{F}_{2},l,s)=1$ and we decompose $\operatorname{Leg}\mathcal{F}_{2}=\mathcal{F}_{2}\boxtimes\mathcal{W}_{d_{2}-1}$ where $\mathcal{F}_{2}$ is a foliation tangent to $\check{s}$ and $\mathcal{W}_{d_{2}-1}$ is a web transverse to $\check{s}$. By the previous argument, $\mathcal{W}_{d_{2}-1}$ and $\operatorname{Leg}\mathcal{F}_{3}$ are transverse to $\mathcal{W}_{d_{1}-\nu_{1}}$ and to each other as well. Thus, by [2, Proposition 3.9], the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{F}_{2}\boxtimes\mathcal{W}_{\nu_{1}})\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-1}\boxtimes\operatorname{Leg}\mathcal{F}_{3})$ is holomorphic along $\check{s}$.
  • If $s\in(\operatorname{Sing}\mathcal{F}_{2}\cap\operatorname{Sing}\mathcal{F}_{3})$ we have $\operatorname{Tang}(\mathcal{F}_{2},l,s)=\operatorname{Tang}(\mathcal{F}_{3},l,s)=1$ and we decompose $\operatorname{Leg}\mathcal{F}_{2}=\mathcal{F}_{2}\boxtimes\mathcal{W}_{d_{2}-1}$ and $\operatorname{Leg}\mathcal{F}_{3}=\mathcal{F}_{3}\boxtimes\mathcal{W}_{d_{3}-1}$ as before. By the previous argument, $\mathcal{W}_{d_{2}-1}$ and $\mathcal{W}_{d_{3}-1}$ are transverse to $\mathcal{W}_{d_{1}-\nu_{1}}$. Thus, by [2, Remark 3.10], the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{F}_{2}\boxtimes\mathcal{F}_{3}\boxtimes\mathcal{W}_{\nu_{1}})\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-1}\boxtimes\mathcal{W}_{d_{3}-1})$ is holomorphic along $\check{s}$.
Now, take $s\in(\Sigma_{\mathcal{F}_{1}}^{rad}\cap\Sigma_{\mathcal{F}_{2}}^{rad})\setminus\Sigma_{\mathcal{F}_{3}}^{rad}$, we have that $s$ is a radial singularity of orders $\nu_{1}-1$ and $\nu_{2}-1$ of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ respectively and we have local decompositions as before: $\operatorname{Leg}\mathcal{F}_{i}=\mathcal{W}_{\nu_{i}}\boxtimes\mathcal{W}_{d_{i}-\nu_{i}}$. We need to analyze two situations:
  • If $s\not\in\operatorname{Sing}\mathcal{F}_{3}$ by Lemma 1.3, we obtain that the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{\nu_{2}})\boxtimes(\operatorname{Leg}\mathcal{F}_{3}\boxtimes\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-\nu_{2}})$ is holomorphic along $\check{s}$
  • if $s\in\operatorname{Sing}\mathcal{F}_{3}$, we can decompose $\operatorname{Leg}\mathcal{F}_{3}=\mathcal{F}_{3}\boxtimes\mathcal{W}_{d_{3}-1}$ with $\mathcal{F}_{3}$ a foliation tangent to $\check{s}$. Again by Lemma 1.4 the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{\nu_{2}}\boxtimes\mathcal{F}_{3})\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-\nu_{2}}\boxtimes\mathcal{W}_{d_{3}-1})$ is holomorphic along $\check{s}$.
In the case where $\check{s}$ is radial singularity of $\mathcal{F}_{1}$, $\mathcal{F}_{2}$ and $\mathcal{F}_{3}$ of respective orders $\nu_{1}-1$, $\nu_{2}-1$, $\nu_{3}-1$ we obtain that $\operatorname{Leg}\mathcal{W}$ is holomorphic along $\check{s}$ by Lemma 1.3.
Now we consider product of two foliations and one invariant line.
@@
ii. If $s\in(\Sigma_{\mathcal{F}_{1}}^{\text{rad}}\cap\Sigma_{\mathcal{F}_{2}}^{\text{rad}})\setminus l$ with radiality orders $\nu_{1}-1$ and $\nu_{2}-1$ respectively. Since $s\not\in l$ we obtain that the direction of $\operatorname{Leg}l$ on the general point of $\check{s}$ does not coincide with any other of $\operatorname{Leg}\mathcal{F}_{1}$ and $\operatorname{Leg}\mathcal{F}_{2}$. So we can write
$\operatorname{Leg}\mathcal{W}=(\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{\nu_{2}})\boxtimes(\operatorname{Leg}l\boxtimes\mathcal{W}_{d_{1}+d_{2}-\nu_{1}-\nu_{2}})$
where $\operatorname{Leg}l\boxtimes\mathcal{W}_{d_{1}+d_{2}-\nu_{1}-\nu_{2}}$ is smooth and transverse to $\check{s}$. Again by Lemma 1.3 we have that $K(\operatorname{Leg}\mathcal{W})$ is holomorphic on $\check{s}$.
iii. If $s\in l\setminus\Sigma_{\mathcal{F}_{2}}^{\text{rad}}$ with $\nu_{1}-1$ being the radiality order of $s$ in $\mathcal{F}_{1}$, then we write $\operatorname{Leg}\mathcal{F}_{1}=\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{d_{1}-\nu_{1}}$. Note that the directions of $\operatorname{Leg}\mathcal{F}_{2}$ and $\mathcal{W}_{d_{1}-\nu_{1}}$ do not coincide at generic $\check{t}\in\check{s}$, and $\operatorname{Tang}(\mathcal{F}_{2},t,s)\leq 1$ (because $s\not\in\Sigma_{\mathcal{F}_{2}}^{\text{rad}}$). Then,
@@
    with $\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-1}$ smooth and transverse to $\check{s}$ and $\mathcal{W}_{1}$ a foliation tangent to $\check{s}$. If $\nu_{1}\geq 2$ the flatness of $\operatorname{Leg}\mathcal{W}$ will follow from [2, Remark 3.10]. On other hand, if $\nu_{1}=1$ we conclude by [4, Theorem 1.1].
iv. If $s\in\Sigma_{\mathcal{F}_{1}}^{\text{rad}}\cap\Sigma_{\mathcal{F}_{2}}^{\text{rad}}\cap l$ with $\nu_{1}-1$ and $\nu_{2}-1$ being the orders of $s$, respectively, in $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, then we can write
$\operatorname{Leg}\mathcal{W}=(\operatorname{Leg}l\boxtimes\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{\nu_{2}})\boxtimes(\mathcal{W}_{d_{1}+d_{2}-\nu_{1}-\nu_{2}})$
with $\mathcal{W}_{d_{1}+d_{2}-\nu_{1}-\nu_{2}}$ smooth and transverse to $\check{s}$. We conclude by Lemma 1.4.
@@
Set $\mathcal{W}^{\prime}=\operatorname{Leg}l_{i}\boxtimes\cdots\boxtimes\operatorname{Leg}l_{k}$ and $\mathcal{W}^{\prime\prime}=\operatorname{Leg}\mathcal{F}_{1}\boxtimes\cdots\boxtimes\operatorname{Leg}\mathcal{F}_{n}$. By [1, lemma 2.1], we have that
By [6, Theorem 4.2], Propositions 2.2, 2.5, and 2.6 we are in condition to apply Corollary 1.6 and conclude that $K(\operatorname{Leg}l_{i}\boxtimes\mathcal{W}^{\prime\prime})=0$. Similarly, we use in addition [1, Theorem 2] to be able to use Corollary 1.7 and obtain $K(\operatorname{Leg}l_{i}\boxtimes\operatorname{Leg}l_{j}\boxtimes\mathcal{W}^{\prime\prime})=0$. Finally, by Proposition 1.5 we have $K(\mathcal{W}^{\prime\prime})=0$ and we finish our proof.
Example 2.7.
Take $\mathcal{F}_{d}$ and $\mathcal{F}_{2d-1}$ Fermat foliations of degree $d$ and $12d-$ 1. By Proposition 2.4, the curve $\operatorname{Tang}(\mathcal{F}_{d},\mathcal{F}_{2d-1})$ is formed by the invariant lines:
[…]
In this way, if $C$ is the reduced curve formed by the product of some of these lines and $\mathcal{W}=C\boxtimes\mathcal{F}_{d}\boxtimes\mathcal{F}_{2d-1}$, we have that $\operatorname{Leg}(\mathcal{W})$ is flat.
@@ 3. Homogeneous Foliations
where $d=d_{1}+\ldots+d_{n}$. So we have that $\operatorname{Leg}\mathcal{W}$ is invariant by the dual maps $\check{h}_{\lambda}(a,b)=(\frac{a}{\lambda},\frac{b}{\lambda})$ and so
@@
Let $\mathcal{H}_{1}$, $\mathcal{H}_{2}$ be convex homogeneous foliations on $\mathbb{P}^{2}$ such that, $\operatorname{Tang}(\mathcal{H}_{1},\mathcal{H}_{2})=L_{\infty}\cup l_{1}\cup\ldots\cup l_{k}$ where $l_{j}$ is an invariant line $\forall j=1,\ldots,k$. Consider the web $\mathcal{W}=\mathcal{H}_{1}\boxtimes\mathcal{H}_{2}$. Then $\operatorname{Leg}\mathcal{W}=\operatorname{Leg}\mathcal{H}_{1}\boxtimes\operatorname{Leg}\mathcal{H}_{2}$ is flat.
@@
If $s\in(\Sigma_{\mathcal{H}_{1}}^{rad}\cup\Sigma_{\mathcal{H}_{2}}^{rad})$, we do a similar analysis to that done in the Proposition 2.2 and obtain that $K(\operatorname{Leg}\mathcal{W})$ is holomorphic over $\check{s}$. Thus $K(\operatorname{Leg}\mathcal{W})$ is holomorphic on ˇ$\check{\mathbb{P}}^{2}_{\mathbb{C}}\setminus\check{O}$ and by Lemma 3.1 we have that $\operatorname{Leg}\mathcal{W}$ is flat.